### v1

@@ Abstract
There is widespread concern about the negative impacts of social media feed ranking algorithms on political polarization. Leveraging advancements in large language models (LLMs), we develop an approach to re-rank feeds in real-time to test the effects of content that is likely to polarize: expressions of antidemocratic attitudes and partisan animosity (AAPA). In a preregistered 10-day field experiment on X/Twitter with 1,256 consented participants, we increase or decrease participants’ exposure to AAPA in their algorithmically curated feeds. We observe more positive outparty feelings when AAPA exposure is decreased and more negative outparty feelings when AAPA exposure is increased. Exposure to AAPA content also results in an immediate increase in negative emotions, such as sadness and anger. The interventions do not significantly impact traditional engagement metrics such as re-post and favorite rates. These findings highlight a potential pathway for developing feed algorithms that mitigate affective polarization by addressing content that undermines the shared values required for a healthy democracy.
keywords:
social media, polarization, feed ranking, field experiment
1 Main
@@
Social media algorithms profoundly impact our lives: they curate what we see [1] in ways that can shape our opinions [2, 3, 4], our moods [5, 6, 7], and our actions [8, 9, 10, 11, 12]. Due to the power that these ranking algorithms have to direct our attention, many have expressed concern that they may be exacerbating negative societal outcomes [13, 14, 15, 16, 17]. Of particular concern is whether feed algorithms cause affective polarization—hostility toward opposing political parties [18, 19, 20]. If social media algorithms are causing affective polarization, they might not only bear responsibility for rising political incivility online [21] but also pose a risk to trust in democratic institutions [22]. In this case, isolating the algorithmic design choices that cause polarization could offer alternative algorithmic approaches.
@@
Research investigating this question has so far yielded mixed results due to limited collaboration between platforms and researchers, frequent changes in feed-ranking algorithms, and the lack of strong incentives for platforms to address the issue [23]. A major hypothesized mechanism for how feed algorithms cause polarization is a self-reinforcing engagement loop: users engage with content aligning with their political views, the feed algorithm interprets this engagement as a positive signal, and the algorithm exposes even more politically aligned content to users, leading to a polarizing cycle. Some studies support this hypothesis, finding that online interactions exacerbate polarization [24], potentially because of the increased visibility of hostile political discussions [25], divisive language [26, 27, 28, 29, 30], and content that reinforces existing beliefs [31]. However, large-scale field experiments aimed at reducing polarization by intervening on the feed algorithm, for example by increasing exposure to out-party content, have found both a decrease [32] and increase [33] in polarization. Similarly, recent large-scale experiments on Facebook and Instagram found no evidence that exposure to in-party sources or a simpler reverse-chronological algorithm affected polarization and political attitudes [34, 35] during the 2020 U.S. election [36]. These mixed results have led to difficulty in identifying what, if any, algorithmic intervention might help.
@@
We can distill the goals of these prior interventions to an alternative hypothesis: that feed algorithms cause affective polarization by exposing us to content that polarizes. An algorithm that upranks genuine political dialogue is less likely to polarize than an algorithm that upranks demagoguery. This hypothesis has been difficult to operationalize into interventions, making studies that intervene on cross-partisan exposure and reverse-chronological ranking attractive but more diffuse in their impact and thus more likely to observe mixed results. However, recent advances in AI large language models (LLMs) now provide an opportunity to create a ranking intervention that more directly targets the focal hypothesis [37]. We draw in particular on a recent large-scale field experiment that articulated eight categories of antidemocratic attitudes and partisan animosity as bipartisan threats to the healthy functioning of democracy [38]. We operationalize these eight categories into an AI classifier that labels expressions of these constructs in social media posts, does so as accurately as trained annotators agree with each other, and produces the desired depolarization effects in a lab setting on a fixed feed [39]. This real-time classification enables us to perform a scalable, content-based, real-time reranking experiment on participants’ own feeds in the field [40].
@@
We conduct a preregistered field experiment on X (Twitter), the most used platform for political discourse in the US [41], that re-ranks participants’ social media content by either increasing or decreasing exposure to content that expresses these eight factors of antidemocratic attitudes and partisan animosity (AAPA). Consented participants (Sec. 6) installed a web extension and experienced this intervention through a real-time reranking of their feeds when they visited the X website over the course of a week. We measured the intervention’s effect on affective polarization [42] and emotional experience [43]. Compared to control conditions that did not rerank the feed, increased AAPA exposure in the ranking algorithm led to colder feelings toward the political outgroup, whereas decreased AAPA exposure led to warmer feelings. These changes also had an effect on participants’ levels of negative emotions—anger and sadness—as measured through in-feed surveys. Our intervention does not lead to statistically significant changes to traditional engagement metrics. However, because we observe higher engagement on political social media content than other content, the engagement-prioritizing nature of current feed algorithms is likely already amplifying AAPA content.
This work demonstrates not just causal evidence of feed algorithm polarization but a potential intervention to mitigate it by downranking content that violates shared democratic values.
2 Exposure to Antidemocratic Attitudes and Partisan Animosity
2.1 Antidemocratic Attitudes and Partisan Animosity (AAPA)
@@
There are many viable theories as to which content, if amplified, could impact affective polarization. In this experiment, we draw specifically on prior work that articulates eight factors as problematic outcomes for healthy democratic functioning, referred to as “antidemocratic attitudes and partisan animosity” [38]: (1) partisan animosity, (2) support for undemocratic practices, (3) support for partisan violence, (4) support for undemocratic candidates, (5) opposition to bipartisanship, (6) social distrust, (7) social distance, and (8) biased evaluation of politicized facts (Sec. A.1). We refer to these eight factors collectively as AAPA (for Antidemocratic Attitudes and Partisan Animosity), and we hypothesize that increased or decreased exposure to AAPA content will cause corresponding changes to affective polarization. Our feed ranking intervention first identifies only political content in the feed, using a broad definition from the Pew Research Center [44] that includes mentions of officials and activists, social issues, news, and current events (Sec. B.3). We then score each political post from zero to eight based on the number of the eight AAPA factors expressed in its text, and we classify a post as AAPA if it reflects at least four of these eight factors. This scoring is performed by a large language model (Sec. B.4), which shows a high correlation with human judgment on identifying these eight factors in social media posts [39].
2.2 Experimental design
@@
Figure 1: Overview of the field experiment on X. (A) Participants complete a pre-experiment survey to rate their feelings toward the opposing party and report their emotions. Then, they are randomly assigned to one of the two parallel experiments, increased or reduced AAPA content exposure, and further randomized into treatment and control groups. After 10 days, they complete a post-experiment survey. (B) Timeline of the experiment, which spans ten days. During the first three days, no intervention is applied to measure the participants’ baseline responses. After day three, participants in the treatment groups received increased or decreased AAPA exposure. In-feed surveys are occasionally shown throughout the baseline and the intervention period. (C) The interventions downranked or upranked AAPA posts in the participants’ feeds, depending on their treatment assignment. In-feed surveys are occasionally added to the feed to capture participants’ in-feed responses.
@@
We designed two parallel experiments to measure the impact of reducing and increasing exposure to AAPA posts (Fig. 1). The first, a “Reduced Exposure” experiment, downranked all AAPA content, pushing it lower in their feeds. In contrast, the second, an “Increased Exposure” experiment, upranked attitudinally-aligned AAPA posts into participants’ feeds. Both experiments are randomized controlled trials. After providing informed consent and installing our web extension, participants were randomly assigned to one of the two experiments and then randomly divided into treatment or control conditions for each experiment. The overall experiment spanned ten days, with the first three days serving as a baseline period, during which no intervention was applied. During the subsequent seven days, participants in the treatment groups received the modified feed (downranked or upranked AAPA posts), while those in the control groups continued to receive the original, unmodified feed.
@@
We assess outcomes at two different time points: post-experiment and in-feed. To measure post-experiment effects, we administered a survey asking participants how warm they feel towards people of the opposing party (0–100 degrees on a feeling thermometer) and about the emotions that they associate with reading social media over the previous week (5-point scale between “never” and “all the time”). We surveyed the participants both before and after the experiment, enabling us to quantify the impact of the intervention at the end of the week-long treatment (Sec. 5.2).
@@
To measure in-feed effects, we added posts containing a small survey directly in the participant’s feed. These surveys, modeled on the ecological momentary assessment approach [45], resemble typical social media posts but are colored to stand out and contain a slider survey instrument. Each survey randomly selects one of two outcomes to assess: (1) affective polarization, where participants rate their feelings toward the opposing party, or (2) emotional responses, where participants report the extent to which they feel two emotions, one randomly selected positive emotion (excited or calm), and one randomly selected negative emotion (angry or sad). The in-feed surveys appeared with a decreasing probability after each feed intervention (post downranking or upranking) and were shown at an equivalent rate and position in the control conditions (Sec. 5.2).
2.3 Feed Reranking
@@
To administer the interventions, we developed a browser extension for Google Chrome and Microsoft Edge that reranks participants’ feeds in real time [40]. The extension activates when participants access the “For You” feed on X. It intercepts the participants’ original feed and sends it to a remote backend, which scores and reranks the posts (Sec. B). This process adds a latency of about three seconds when opening the feed for the first time, which is consistent across all experimental conditions. To rerank the original feeds, we first identify political posts (Sec. B.3) and then determine whether those posts are classified as AAPA.
@@
In the Reduced Exposure treatment condition, we downrank all AAPA posts in the participants’ feeds. If participants continue to scroll down far enough, e.g., across several incremental loads of more content, these demoted posts will appear (Fig. 9). Before downranking all the AAPA posts, we randomly select the position of one AAPA post and, if the in-feed survey is sampled to be included, insert the post with the survey in the next available position (Sec. A.2.1). This survey position is consistent across treatment and control conditions. Thus, participants in the control condition will see AAPA posts and the in-feed survey, while participants in the treatment condition will see the survey in the position where the first AAPA post used to be, but all AAPA posts are placed further down. The survey is limited to only a few times per day to avoid fatigue (Sec. 5.2). In the Increased Exposure experiment, since we lack access to the platform’s full post inventory, we sourced AAPA posts from the inventory of all posts in the “For You” feeds of all other study participants. We selected a random position in the original feed and upranked a random AAPA post to that position. The post containing the survey, if included, is inserted in the next available position. In the control condition, the survey is still added in the randomly sampled position, but no post is upranked. As in the Reduced Exposure experiment, the survey is placed consistently in the treatment and control conditions. During continuous scrolling sessions, this process repeats approximately every 30 posts (Sec. A.2.2).
The Reduced Exposure treatment downranked 85.4 AAPA posts on average (median = 22) per user per day, and the Increased Exposure upranked 10.9 AAPA posts on average (median = 5.43) per user per day.
3 Results
@@
We recruited participants from two online survey platforms, CloudResearch and Bovitz-Forthright [46]. We targeted participants who use X, are 18 or older, reside in the United States, and self-identify as either Democrats or Republicans (Sec. C). After obtaining informed consent, participants installed our web extension. To ensure the participants had enough political content, we used a pre-screening task to filter for users whose “For You” feeds contained at least 5% of posts related to politics or social issues. The participants were recruited during the U.S. election season, between July 7th and August 14th, 2024, and instructed to use X daily only from the web interface.
@@
We preregistered a target of at least 1,100 participants to complete the full study, broken down as 500 in the Increased Exposure experiment and 600 in the Reduced Exposure experiment based on power analyses performed on pilot studies. Toward this goal, a total of 1,662 participants completed the initial survey and qualified for inclusion in the study (Fig. 13). Participants were randomly assigned to either the Increased or Reduced Exposure group until the preregistered experiment-specific quotas were reached [47]. Of the qualified participants, N=1,256 completed the full study by responding to follow-up surveys the day after the end of the study: 727 in the Reduced Exposure experiment and 529 in the Increased Exposure experiment. Both treatment and control groups were well balanced across covariates, and there was no evidence of selective attrition between conditions (Sec. 5.3).
@@
On average, participants viewed 158 posts per day (median = 94) and spent 43.8 minutes (median = 19.2 minutes) per day with a browser tab in the foreground on their X feed. During the baseline period, an average of 32% of the posts participants viewed were related to politics or social issues. Of these, 33.1% were classified as AAPA, which comprised an average of 10.6% of the total feed content.
@@
The interventions had the intended effects (Fig. 11). During the baseline period, participants across all groups were exposed to political posts with similar average AAPA scores. However, after the intervention began, participants in the treatment condition of the Reduced Exposure experiment experienced a noticeable decrease in exposure to such posts (-91.2% compared to control), while those in the treatment condition of the Increased Exposure experiment were exposed to more AAPA posts (+23.1% compared to control), resulting in higher average scores.
@@
Over the full ten days of the experiment, participants answered an average of 13.8 (median = 11) in-feed affective polarization questions in the Increased Exposure experiment and 12.6 (median = 9) in the Reduced Exposure experiment. Similarly, they answered an average of 14.0 (median = 11) emotion surveys in the Increased Exposure experiment and 12.5 (median = 9) in the Reduced Exposure experiment–each survey requiring ratings for two emotions.
@@
In the post-survey, we asked participants whether the browser extension had impacted their experience on X. Across conditions, most participants reported that they did not notice any impact in their online experience (74.2%), while others guessed changes that were not part of the intervention (Sec. E.6). These results indicate a limited understanding of the experimental feed ranking intervention and suggest that it is less likely that demand effects might be influencing the outcomes.
@@
3.1 Affective polarization
Exposure to AAPA significantly influenced affective polarization (Fig. 2). Participants in the reduced AAPA exposure experiment reported a significant post-experiment effect, corresponding to an outparty warmth increase of 2.11 degrees (95% CIs: [0.15, 4.06]; P=0.035). Conversely, participants in the Increased Exposure experiment exhibited a symmetric reduction in affective warmth toward the opposing party of -2.48 degrees (95% CIs: [-4.79, -0.17]; P=0.036). We did not observe any significant heterogeneity in treatment effects across the preregistered moderators (Sec. E.5), including party identification, indicating that the effect was consistent and bipartisan. As is common [48], we contextualize these effects against the rising levels of partisan animosity in American society over time: our effects correspond to a decrease and increase in affective polarization equivalent to 3 and 3.6 years, respectively. We observed even stronger effects with the in-feed assessment, with Reduced Exposure causing an increase of 3.24 degrees (95% CIs: [1.21, 5.27]; P=0.002) and Increased Exposure causing a cooling of -2.56 degrees (95% CIs: [-4.53, -0.59]; P=0.011).
@@ 3.1 Affective polarization
In an exploratory analysis, we investigated which AAPA factors were predictive of changes in affective polarization. Specifically, a higher proportion of posts supporting Partisan Animosity (95% CIs: [-13.49, -2.61]; P=0.003), Opposition to Bipartisan Cooperation (95% CIs: [-26.55, -4.40]; P=0.006), and Biased Evaluation of Politicized Facts (95% CIs: [-12.83, -2.21]; P=0.005;) predicted a colder feeling toward the opposing party by the end of the study (Fig. 17).
@@
Figure 3: Effects of reducing and increasing exposure to AAPA content in participants’ feeds on their experiences of emotion (in-feed scale ranged from 0: “none at all” to 100: “extremely”, post-experiment scale ranged from 1: “never” to 5: “all the time”) relative to the corresponding control group. Participants were surveyed within the feed during the intervention and off-platform after the experiment. The filled circles represent statistical significance ($P<0.05$, adjusted for multiple hypothesis testing), and the error bars represent 95% confidence intervals.
3.2 Emotions

### v2

@@ Abstract
Today, social media platforms hold sole power to study the effects of feed ranking algorithms. We developed a platform-independent method that reranks participants’ feeds in real-time and used this method to conduct a preregistered 10-day field experiment with 1,256 participants on X during the 2024 U.S. presidential campaign. Our experiment used a large language model to rerank posts that expressed antidemocratic attitudes and partisan animosity (AAPA). Decreasing or increasing AAPA exposure shifted out-party partisan animosity by two points on a 100-point feeling thermometer, with no detectable differences across party lines, providing causal evidence that exposure to AAPA content alters affective polarization. This work establishes a method to study feed algorithms without requiring platform cooperation, enabling independent evaluation of ranking interventions in naturalistic settings.
Social media algorithms profoundly impact our lives: they curate what we see  (?) in ways that can shape our opinions  (?, ?, ?), our moods  (?, ?, ?), and our actions  (?, ?, ?, ?, ?). Due to the power that these ranking algorithms have to direct our attention, the research literature has articulated many theories and results detailing the impact that ranking algorithms have on us  (?, ?, ?, ?, ?). However, validating these theories and results has remained extremely difficult because the ranking algorithm behavior is determined by the social media platforms, and only the platforms themselves can test alternative feed designs and causally assess their impact. Platforms, however, face political and financial pressures that constrain the kinds of experiments they can launch and share  (?). Concerns about lawsuits and the need to preserve engagement-driven revenue further limit what platforms are willing to test, leaving massive gaps in the design space of ranking algorithms that have been explored in naturalistic settings and at scale.
To address this gap, in this work, we present an approach that enables researchers to rerank participants’ social media feeds in real-time as they browse, without requiring platform permission or cooperation. We built a browser extension—a small add-on to a web browser that modifies how web pages appear or behave, similar to an ad blocker. Our extension intercepts and modifies X’s web-based feed in real-time and reranks the feed using Large Language Model (LLM)-based rescoring, with only a negligible increase in page load time. This web extension allows us to rerank content according to experimentally controlled conditions. The design opens a new paradigm for algorithmic experimentation: it provides external researchers with a tool for conducting independent field experiments and evaluating the causal effects of algorithmic content curation on user attitudes and behaviors while preserving ecological validity.
@@
This capability allowed us to investigate a pressing question: can feed algorithms cause affective polarization—hostility toward opposing political parties  (?, ?, ?, ?)? This concern has grown since the 2016 U.S. presidential election  (?), and the debate remains ongoing after the 2020 and 2024 elections. If social media algorithms are causing affective polarization, they might not only bear responsibility for rising political incivility online  (?) but also pose a risk to trust in democratic institutions  (?). In this case, isolating the algorithmic design choices that cause polarization could offer alternative algorithmic approaches  (?).
@@
A major hypothesized mechanism for how feed algorithms cause polarization is a self-reinforcing engagement loop: users engage with content aligning with their political views, the feed algorithm interprets this engagement as a positive signal, and the algorithm exposes even more politically aligned content to users, leading to a polarizing cycle. Some studies support this hypothesis, finding that online interactions exacerbate polarization  (?), potentially because of the increased visibility of hostile political discussions  (?), divisive language  (?, ?, ?, ?, ?), and content that reinforces existing beliefs  (?). However, large-scale field experiments aimed at reducing polarization by intervening on the feed algorithm—for example, by increasing exposure to out-party content—have found both a decrease  (?) and an increase  (?) in polarization. Similarly, recent large-scale experiments on Facebook and Instagram found no evidence that reduced exposure to in-party sources or a simpler reverse-chronological algorithm affected polarization and political attitudes  (?, ?) during the 2020 U.S. election. These mixed results reveal the difficulty in identifying what, if any, algorithmic intervention might help reduce polarization, especially during politically charged times.
@@
We distill the goals of these prior interventions to a direct hypothesis that we can operationalize via real-time LLM reranking: that feed algorithms cause affective polarization by exposing users specifically to content that polarizes. An algorithm that upranks content reflecting genuine political dialogue is less likely to polarize than one that upranks demagoguery. This content-focused hypothesis has been difficult to operationalize into interventions, making studies that intervene on cross-partisan exposure and reverse-chronological ranking attractive but more diffuse in their impact and thus more likely to observe mixed results. However, by connecting our real-time reranking infrastructure with recent advances in LLMs, we can create a ranking intervention that more directly targets the focal hypothesis  (?) without needing platform collaboration. We draw, in particular, on a recent large-scale field experiment that articulated eight categories of antidemocratic attitudes and partisan animosity as bipartisan threats to the healthy functioning of democracy  (?). We operationalize these eight categories into an AI classifier that labels expressions of these constructs in social media posts, does so with accuracy comparable to trained annotators, and produces depolarization effects in a lab setting on a fixed feed  (?). This real-time classification enables us to perform a scalable, content-based, reranking experiment on participants’ own feeds in the field  (?).
@@
We conducted a preregistered field experiment on X, the most used social media platform for political discourse in the U.S.  (?), using our extension to dynamically rerank participants’ social media content by either increasing or decreasing exposure to content that expresses these eight factors of antidemocratic attitudes and partisan animosity (AAPA) over the course of a week. The experiment was conducted during a pivotal moment in the 2024 United States election cycle, from July to August 2024—an important period for understanding how social media feeds impact affective polarization. Major political events during the study period included the attempted assassination of Donald Trump, the withdrawal of Joe Biden from the 2024 presidential race, and the nomination of Kamala Harris as the Democratic Party’s candidate. These events allow us to examine the impact of heterogeneous AAPA content on partisan polarization and hostility. We measured the intervention’s effect on affective polarization  (?) and emotional experience  (?). Compared to control conditions that did not rerank the feed, decreased AAPA exposure led to warmer feelings toward the political outgroup, whereas increased AAPA exposure led to colder feelings. These changes also affected participants’ levels of negative emotions (anger and sadness) as measured through in-feed surveys.
Exposure to Antidemocratic Attitudes and Partisan Animosity
Antidemocratic Attitudes and Partisan Animosity (AAPA)
@@
There are many viable theories as to which content, if amplified, could impact affective polarization. In this experiment, we draw specifically on prior work that articulates eight factors as problematic outcomes for healthy democratic functioning, referred to as “antidemocratic attitudes and partisan animosity”  (?): (1) partisan animosity, (2) support for undemocratic practices, (3) support for partisan violence, (4) support for undemocratic candidates, (5) opposition to bipartisanship, (6) social distrust, (7) social distance, and (8) biased evaluation of politicized facts (Appendix S3.1). We refer to these eight factors collectively as AAPA (for Antidemocratic Attitudes and Partisan Animosity), and hypothesize that increased or decreased exposure to AAPA content causes corresponding changes in affective polarization. Our feed ranking intervention first identified only political content in the feed, using a broad definition from the Pew Research Center  (?) that includes mentions of officials and activists, social issues, news, and current events (Appendix S4.3). We then scored each political post from zero to eight based on the number of the eight AAPA factors expressed in its text, and we classified a post as AAPA if it reflected at least four of these eight factors. This scoring was performed by a large language model (Appendix S4.4), which showed a high correlation with human judgment on identifying these eight factors in social media posts  (?).
Experimental Design
@@
Figure 1: Overview of the field experiment on X. (A) Participants completed a pre-experiment survey to rate their feelings toward the opposing party and report their emotions. Then, they were randomly assigned to one of the two parallel experiments, reduced or increased AAPA content exposure, and further randomized into treatment and control groups. After ten days, they completed a post-experiment survey. (B) Timeline of the experiment. During the first three days, no intervention was applied to measure the participants’ baseline responses. For the next seven days, participants in the treatment groups received increased or decreased AAPA exposure. In-feed surveys were periodically shown throughout the baseline and the intervention period. (C) The interventions downranked or upranked AAPA posts in the participants’ feeds, depending on their treatment assignment.
@@
We designed two parallel experiments to measure the impact of reducing and increasing exposure to AAPA posts (Fig. 1). The first, a “Reduced Exposure” experiment, downranked all AAPA content, pushing it lower in their feeds. In contrast, the second, an “Increased Exposure” experiment upranked AAPA posts into participants’ feeds. These posts came from the inventory of posts recommended by X to all participants with the same political leaning. Both experiments were randomized controlled trials. After providing informed consent and installing our web extension, participants were randomly assigned to one of the two experiments and then randomly divided into treatment or control conditions for each experiment. The overall experiment spanned ten days, with the first three days serving as a baseline period, during which no intervention was applied. During the subsequent seven days, participants in the treatment groups received the modified feed (downranked or upranked AAPA posts), while those in the control groups continued to receive the original, unmodified feed.
@@
We assessed outcomes at two different time points: post-experiment and in-feed. To measure post-experiment effects, we administered a survey asking participants how warmly they feel towards members of the opposing party (0–100 degrees on a feeling thermometer) and about the various emotions (e.g., anger, sadness, excitement, calm) they felt when reading social media over the previous week (5-point scale between “never” and “all the time”). We surveyed the participants both before and after the experiment, enabling us to quantify the impact of the intervention at the end of the week-long treatment (Appendix S1.2).
@@
To measure in-feed effects, we added posts containing a small survey directly in the participant’s feed. These surveys, modeled on ecological momentary assessments  (?), resembled typical social media posts but were colored to stand out and contained a slider survey instrument. Each survey randomly selected one of two outcomes to assess: (1) affective polarization, where participants rated their feelings toward the opposing party, or (2) emotional experience, where participants reported the extent to which they felt two emotions: one randomly selected positive emotion (excited or calm), and one randomly selected negative emotion (angry or sad). For both experiments, the in-feed surveys were shown at an equivalent rate and position in the control conditions (Appendix S1.2). To reduce fatigue effects, each time participants completed the in-feed survey, they were less likely to receive another one on the same day.
Feed Reranking
@@
To administer the interventions, we developed a browser extension for Google Chrome and Microsoft Edge that reranked participants’ feeds in real-time  (?). The extension activated when participants accessed the “For you” feed on X. It intercepted the participants’ original feed and sent it to a remote backend, which scored and reranked the posts (Appendix S4). This process added a latency of about three seconds when opening the feed for the first time, which was consistent across all experimental conditions. To rerank the original feeds, we first identified political posts (Appendix S4.3), then classified whether and how much AAPA the political posts contained.
@@
In the Reduced Exposure treatment condition, we downranked all AAPA posts in the participants’ feeds. If participants continued to scroll down far enough, e.g., across several incremental loads of more content, these demoted posts would appear (Fig. S12). Before downranking all the AAPA posts, we randomly selected the position of one AAPA post and, if the in-feed survey was sampled to be included, inserted the post with the survey in the next available position (Appendix S3.2.1). This survey position was consistent across treatment and control conditions. Thus, participants in the control condition saw AAPA posts and the in-feed survey, while participants in the treatment condition saw the survey in the position where the first AAPA post used to be, but all AAPA posts were placed further down. The survey was limited to only a few times per day to avoid participant fatigue (Appendix S1.2). In the Increased Exposure experiment, since we lacked access to the platform’s full post inventory, we sourced AAPA posts from the inventory of all posts in the “For you” feeds of all other study participants. We selected a random position in the original feed and upranked a random AAPA post to that position. The post containing the survey, if included, was inserted in the next available position. In the control condition, the survey was still added in the randomly sampled position, but no post was upranked. As in the Reduced Exposure experiment, the survey was placed consistently in the treatment and control conditions. During continuous scrolling sessions, this process repeated approximately every 30 posts (Appendix S3.2.2).
The Reduced Exposure treatment downranked 85.4 AAPA posts on average (median = 22) per user per day, and the Increased Exposure upranked 10.9 AAPA posts on average (median = 5.43) per user per day (Appendix S3.2.3). To avoid excessive exposure to polarizing content, we upranked fewer but higher-scoring AAPA posts. Despite the difference in the number of posts affected by the two interventions, the average AAPA score of the feed consumed during treatment days exhibited symmetric changes across the two experiments (Fig. S11).
Results
@@
We recruited participants from two online survey platforms, CloudResearch and Bovitz  (?). We targeted participants who used X, were 18 or older, resided in the United States, and self-identified as either Democrats or Republicans (Appendix S5). After giving informed consent, participants installed our web extension. To ensure the participants had enough political content, we used a pre-screening task to filter for users whose “For you” feeds contained at least 5% of posts related to politics or social issues, retaining 73.4% of the screened participants. The participants were recruited during the U.S. election season, between July 7th and August 14th, 2024, and instructed to use X daily only from the web interface.
@@
We preregistered a target of at least 1,100 participants to complete the full study (600 in the Reduced Exposure experiment; 500 in the Increased Exposure experiment) based on power analyses performed on pilot studies  (?). Toward this goal, a total of 1,662 participants completed the initial survey and qualified for inclusion in the study. Participants were randomly assigned to either the Reduced or Increased Exposure experiment until the preregistered experiment-specific quotas were reached. Of the qualified participants, N=1,256 completed the full study by responding to a follow-up survey at the end of the experiment: 727 in the Reduced Exposure experiment and 529 in the Increased Exposure experiment (Fig. S1 shows the participant recruitment funnel). Both treatment and control groups were well balanced across covariates, and there was no evidence of differential attrition between conditions (Appendix S1.3). To maintain a consistent sample across the post-experiment and in-feed analyses, the reported treatment effects are based on participants who completed both the post-experiment survey and at least one in-feed survey for the corresponding outcome (N = 1,090). The results are similar when the full population is considered (Appendix S11.1 and S11.2).
Participants received \$20 upon completing the study. The sample included 48% women, 50% men, and 2% who identified as another gender. Participants ranged in age from 18 to 81 years (mean = 41.6, SD = 14.0). Party affiliation was 66% Democrats and 34% Republicans. Educational attainment was as follows: 17 (1%) had less than a high school degree, 158 (13%) were high school graduates, 267 (21%) had some college but no degree, 145 (12%) held an associate degree, 419 (33%) held a bachelor’s degree, 189 (15%) held a master’s degree, 30 (2%) held a professional degree (e.g., JD, MD), and 31 (2%) held a doctoral degree. Participants reported their race/ethnicity, and were allowed to choose multiple responses. Overall, 945 (75%) identified as White, 183 (15%) as Black or African American, 153 (12%) as Hispanic or Latino, 66 (5%) as Asian, 34 (3%) as South Asian, 28 (2%) as American Indian or Alaska Native, 5 (<<1%) as Middle Eastern, and 9 (<<1%) as Other.
@@
On average, participants viewed 154 posts per day (median = 92) and spent 24.1 minutes (median = 13.1 minutes) per day with a browser tab in the foreground on their X feed. During the baseline period, an average of 31.6% of the posts participants viewed were related to politics or social issues. Of these posts, 32% were classified as AAPA, which comprised an average of 10.1% of the total feed content.
@@
The interventions had the intended effects (Fig. S9). During the baseline period, participants across all experimental conditions were exposed to a similar fraction of AAPA posts. However, after the intervention began, participants in the treatment condition of the Reduced Exposure experiment experienced a noticeable decrease in exposure to such posts (treatment: 1.04%, control: 10% AAPA posts of all feed exposures; 89.6% relative reduction), while those in the treatment condition of the Increased Exposure experiment were exposed to more AAPA posts (treatment: 13.4%, control: 10.6% AAPA posts; 26.4% relative increase).
@@
Over the full ten days of the experiment, participants answered an average of 13 (median = 9) in-feed affective polarization questions in the Reduced Exposure experiment and 14.7 (median = 11) in the Increased Exposure experiment. Similarly, they answered an average of 13.2 (median = 9) in-feed emotion surveys in the Reduced Exposure experiment and 15.0 (median = 12) in the Increased Exposure experiment, with each survey asking them to rate two emotions.
@@
In the post-survey, we asked participants whether the browser extension had impacted their experience on X. Across conditions, most participants reported that they did not notice any impact on their online experience (74.2%), while others guessed changes that were not part of the intervention (Appendix S11.4). These results indicate a limited awareness of the experimental feed ranking intervention, reducing the likelihood that the observed results were due to demand effects.
Affective Polarization
Exposure to AAPA significantly influenced affective polarization (Fig. 2). Participants in the reduced AAPA exposure experiment reported a significant post-experiment effect, corresponding to an outparty warmth increase of 2.11 degrees (95% CIs: [0.15, 4.06]; P=0.035). Conversely, participants in the Increased Exposure experiment exhibited a symmetric reduction in affective warmth toward the opposing party of -2.48 degrees (95% CIs: [-4.79, -0.17]; P=0.036). We did not observe any significant heterogeneity in treatment effects across the preregistered moderators (Appendix S8), including party identification, indicating that the effect was consistent and bipartisan. Like other studies  (?), we contextualize these effects against the rising levels of partisan animosity in American society over time: our effects correspond to a decrease and increase in affective polarization equivalent to 3 and 3.6 years, respectively. We observed even stronger effects with the in-feed assessment, with Reduced Exposure causing an increase of 3.24 degrees (95% CIs: [1.211.215.275.27]; P=0.002) and Increased Exposure causing a cooling of -2.56 degrees (95% CIs: [-4.53, -0.59]; P=0.011). Supplementary exploratory analysis reweighting our sample to match the population of X users reduced our statistical power but replicated our results, indicating substantial robustness (Appendix S10). The only exception was the post-treatment effects in the Reduced Exposure experiment, which had a similar point estimate but were not statistically significant.
@@
In an exploratory analysis, we investigated which types of AAPA content were predictive of changes in affective polarization. Specifically, a higher proportion of posts supporting Partisan Animosity (95% CIs: [-13.49, -2.61]; P=0.003), Opposition to Bipartisan Cooperation (95% CIs: [-26.55, -4.40]; P=0.006), and Biased Evaluation of Politicized Facts (95% CIs: [-12.83, -2.21]; P=0.005;) predicted a colder feeling toward the opposing party by the end of the study (Fig. S22).
In addition to affective polarization (or partisan animosity), we also surveyed participants about their attitudes toward the remaining seven AAPA factors. We found that, although reduced and increased exposure to AAPA content impacted affective polarization, it had no significant effects on these other outcomes (Appendix S7). These results are consistent with recent studies finding that changes in affective polarization are not necessarily linked to changes in antidemocratic attitudes  (?, ?).
@@
Figure 3: Effects of reducing and increasing exposure to AAPA content in participants’ feeds on their experiences of emotion, relative to the corresponding control group. Participants were surveyed within the feed during the intervention (scale ranged from 0: “none at all” to 100: “extremely”) and off-platform after the experiment (scale ranged from 1: “never” to 5: “all the time”). The filled circles represent statistical significance ($P<0.05$, adjusted for multiple hypothesis testing), and the error bars represent 95% confidence intervals.
Emotions